### v2

# Adaptive Spatial Goodness Encoding: Advancing and Scaling Forward-Forward Learning

Abstract
The Forward-Forward (FF) algorithm offers a promising alternative to backpropagation (BP). Despite advancements in recent FF-based extensions, which have enhanced the original algorithm and adapted it to convolutional neural networks (CNNs), they often suffer from limited representational capacity and poor scalability to large-scale datasets, primarily due to exploding channel dimensionality. In this work, we propose adaptive spatial goodness encoding (ASGE), a new FF-based training framework tailored for CNNs. ASGE leverages feature maps to compute spatially-aware goodness representations at each layer, enabling layer-wise supervision. Crucially, this approach decouples classification complexity from channel dimensionality, thereby addressing the issue of channel explosion and achieving competitive performance compared to other BP alternatives. ASGE outperforms all other FF-based approaches across multiple benchmarks, delivering test accuracies of 99.65% on MNIST, 93.41% on FashionMNIST, 90.62% on CIFAR-10, and 65.42% on CIFAR-100. Moreover, we present the first successful application of FF-based training to ImageNet, with Top-1 and Top-5 accuracies of 51.58% and 75.23%. Furthermore, we propose three prediction strategies to achieve flexible trade-offs among accuracy, parameters and memory usage, enabling deployment under diverse resource constraints.
Codehttps://github.com/KAI-Laboratory/ASGE
Introduction
Artificial Neural Networks (ANNs) have become a cornerstone of modern machine learning, largely owing to the effectiveness of the backpropagation (BP) algorithm (Rumelhart et al. 1986). BP has driven the success of ANNs across a diverse range of applications (Goodfellow et al. 2014; He et al. 2015; LeCun et al. 2015; Kingma and Ba 2017; Vaswani et al. 2023) by enabling efficient training through gradient-based optimization. The algorithm proceeds in two phases: a forward pass computes the network’s predictions and their associated error relative to the ground truth, followed by a backward pass that propagates this error through the network. This yields gradients for each layer’s parameters (weights and biases), which are then used by optimization methods such as stochastic gradient descent (SGD) (Bottou 2010) to iteratively minimize the prediction error. However, BP suffers from inherent limitations. It requires symmetric weight matrices for the forward and backward passes to compute parameter gradients via the chain rule (Grossberg 1987). Additionally, it necessitates the storage of intermediate activations until parameter updates occur (Czarnecki et al. 2017). These constraints incur substantial computational and memory overhead, while simultaneously restricting the degree of parallelization, ultimately rendering BP inefficient and ill-suited for deployment on resource-constrained hardware.
In response, numerous backpropagation-free (BP-free) strategies have emerged to address these limitations, including Target Propagation (TP) (Bengio 2014), Feedback Alignment (FA) (Lillicrap et al. 2016), Direct Feedback Alignment (DFA) (Nøkland 2016), Deep Random Target Projection (DRTP) (Frenkel et al. 2021), and Recursive Local Representation Alignment (rLRA) (Ororbia et al. 2023), alongside other biologically inspired or local learning methods (Hebb 2002; Moraitis et al. 2022; Dellaferrera and Kreiman 2023). Although these approaches alleviate some of the BP limitations, they continue to face challenges in balancing the trade-offs between prediction accuracy, memory efficiency and computational cost.
Figure 1: Layer-wise training accuracy (layers 2-7) of VGG8 trained with $F$ on CIFAR-10. First and final layers excluded. Left: positive samples. Right: negative samples.
A recent promising alternative is the Forward-Forward algorithm (FF) (Hinton 2022), which employs a layer-wise training strategy through dual forward passes: one with positive samples and the other with negative samples. Each layer independently updates its parameters to maximize a “goodness” metric for positive inputs while minimizing it for negative ones. By eliminating the need for backward passes, FF significantly reduces memory overhead and facilitates pipe-lined parallelism. However, Figure 1 illustrates a critical limitation of FF when applied to convolutional neural networks (CNNs): prediction accuracy tends to deteriorate progressively with network depth. This indicates a weakening capacity for feature abstraction in deeper layers, thereby constraining the algorithm’s effectiveness in deep architectures and on complex tasks. Recent enhancements, such as Channel-wise Competitive (CwC) learning (Papachristodoulou et al. 2024) and DeeperForward (Sun et al. 2025), attempt to mitigate this issue. While partially effective, they introduce an undesirable coupling between channel dimensionality and classification complexity. Specifically, they require layer-wise channels $C_{l}$ to scale linearly with classes $N$, i.e., $C_{l}\in\Theta(N)$. This scaling becomes impractical for large-scale datasets like ImageNet (Russakovsky et al. 2015) due to the dramatic increase in channel dimensionality.
To address both the representational degradation and scalability bottleneck in existing FF-based methods, we propose an enhanced FF-based training framework that simultaneously improves representational capacity and scalability to large-scale datasets. Our key contributions are as follows:
  • Adaptive Spatial Goodness Encoding (ASGE) Framework: A novel BP-free training method that decouples channel dimensionality from classification complexity, enforcing constant channel scaling ($C_{l}\in\Theta(1)$). This enables efficient and scalable optimization on large-category datasets.
  • ImageNet Breakthrough: Comprehensive experiments across various image classification tasks, and first successful scaling of FF-based methods to ImageNet.
Background
BP (Rumelhart et al. 1986) enables end-to-end optimization by propagating errors from the output layer back through the network. However, it suffers from several inherent limitations. First, it requires weight symmetry between the forward and backward passes (Grossberg 1987). Second, BP introduces the update locking problem (Czarnecki et al. 2017), wherein weight updates cannot be performed until the entire forward pass is completed. These constraints result in high memory overhead due to the need to store intermediate activations and hinder layer-wise parallelism, ultimately rendering BP inefficient for deployment on resource-constrained hardware.
To overcome the limitations of backpropagation, numerous alternatives have been proposed that either avoid or approximate its core mechanism. Feedback Alignment (FA) (Lillicrap et al. 2016) eliminates the requirement for weight symmetry in the backward pass by using fixed, randomly initialized feedback weights. DFA (Nøkland 2016) extends this by propagating output layer errors to hidden layers, bypassing the backward pathway altogether. Another related method, rLRA (Ororbia et al. 2023) enables weight updates without computing global gradients by recursively aligning each layer’s output with a synthetic target derived from the layer above. PEPITA (Dellaferrera and Kreiman 2023) replaces the backward pass with a second forward pass, in which the input is modulated according to the output-layer errors from the first forward pass. Notably, these approaches still necessitate full forward passes to compute global errors or store intermediate activations prior to updating weights. To eliminate update locking, local learning rules—which allow layers to update independently without waiting for the entire forward pass—are desirable. DRTP (Frenkel et al. 2021) achieves this by directly projecting true labels to each layer, generating layer-wise error signals that support local updates without waiting for the full forward pass. Another approach is Hebbian learning (Hebb 2002), a biologically inspired local learning rule that strengthens synaptic connections when pre-synaptic and post-synaptic neurons are activated simultaneously. Despite its biological plausibility, Hebbian learning does not support supervised learning or error correction. SoftHebb (Moraitis et al. 2022; Journé et al. 2023) extends Hebbian learning by incorporating soft winner-take-all dynamics and differentiable surrogate gradients, enabling the training of deeper networks while preserving the characteristics of local learning. Similarly, Predictive Coding (Rao and Ballard 1999; Millidge et al. 2022) offers a local learning framework in which each layer predicts the activity of the layer below, using the resulting prediction error to locally update both prediction and representation neurons.
The FF algorithm (Hinton 2022) introduces a fundamentally different training paradigm that entirely avoids the need for backpropagation. It employs a layer-wise learning mechanism in which each layer is independently supervised through two distinct forward passes—one with positive data and the other with negative data. Each layer computes a “goodness” score, defined as the sum of squared activations, and is trained using a binary cross-entropy loss to classify inputs as positive if the goodness exceeds a predefined threshold, and negative otherwise. Several extensions have been proposed to enhance the performance of FF. Contrastive Forward-Forward (Aghagolzadeh and Ezoji 2025) redefines goodness as the dot-product similarity between representations of different samples and introduces layer-wise supervised contrastive learning to improve the discriminative power of the network. SymBa (Lee and Song 2023) addresses the inherent asymmetry in binary cross-entropy loss by introducing a symmetric loss function with respect to positive and negative data, leading to improved convergence stability. SFFA (Terres-Escudero et al. 2025) further enforces symmetry by dividing each layer’s neurons into positive and negative subsets, selectively activating them based on the input’s polarity. To overcome the lack of inter-layer interaction in FF, Layer Collaboration (Lorberbom et al. 2023) proposes a mechanism that allows goodness information to flow across layers, improving representational coherence and overall accuracy. However, these variants have not yet been extended to CNNs, limiting their applicability to more complex classification tasks.
Several recent efforts have explored adapting the FF paradigm to CNNs (Scodellaro et al. 2024; Chen et al. 2025) by injecting polarity information into the global image space to generate positive and negative samples instead of using one-hot label embeddings as in the original FF. Other approaches take this further by eliminating the reliance on explicitly generated negative samples. For example, Cascaded Forward (CaFo) (Zhao et al. 2023) first trains the convolutional backbone using DFA, followed by training a set of layer-specific predictors, both stages relying solely on the original data. However, it suffers from slow convergence. A more recent state-of-the-art method is CwC learning (Papachristodoulou et al. 2024), which partitions the output channels of each layer based on the number of target classes. It calculates a group-wise goodness score for each class-aligned channel group and then uses these scores for classification (Figure 2). CwC has been shown to progressively enhance class-specific feature activation in deeper layers, partially addressing the representation degradation (Figure 1). DeeperForward (Sun et al. 2025) further extends CwC to deeper convolutional architectures and introduces a parallel training strategy to enhance efficiency. However, a significant limitation of CwC lies in its tight coupling between the number of channels and the number of classes. This design causes the channel dimensionality to scale linearly with the number of categories, severely limiting their performance on large-scale benchmarks and rendering them impractical for scaling to ImageNet. In light of these challenges, our work aims to further optimize FF-based training approaches for CNNs, with the goal of achieving improved scalability and performance on large-scale classification tasks.
Method
Figure 2: (a) Top: ASGE architecture with prediction strategies (Fusion Pred: Cumulative Layer Fusion Prediction, Best Pred: Best Layer Selection Prediction, Last Pred: Last Layer Only Prediction); Bottom: Layer-wise adaptive spatial goodness. (b) CwC goodness.
This section presents a backpropagation-free training framework built upon the proposed ASGE, a new mechanism that enhances representational capacity and overcomes the scalability limitations of prior FF-based methods in large-scale image classification tasks. ASGE trains each convolutional layer independently through channel-adaptive spatial goodness supervision, completely eliminating backward passes. We first introduce the core design of the ASGE mechanism, then detail the overall training framework.
Adaptive Spatial Goodness Encoding
Spatial Goodness Extraction
Figure 2 illustrates the fundamental difference in goodness vector construction between our approach and CwC (Papachristodoulou et al. 2024). We propose a spatially grounded, class-agnostic method, where spatial goodness is defined as the energy within localized regions of each channel. This design explicitly decouples goodness computation from class count, enabling flexible supervision while ensuring scalability to large classification tasks. Formally, the feature maps of a given layer $l$ are denoted as $\mathbf{Y}_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ (post-activation), where $C_{l}$ represents channels and $H_{l}\times W_{l}$ spatial dimensions. To standardize the quantification of localized regions, each $H_{l}\times W_{l}$ feature map is evenly partitioned into $P_{l}\times P_{l}$ patches. Accordingly, $\mathbf{Y}_{l}$ is reshaped into:
$\mathbf{\tilde{Y}}_{l}\in\mathbb{R}^{C_{l}\times P_{l}\times P_{l}\times H^{{}^{\prime}}_{l}\times W^{{}^{\prime}}_{l}}$ (1)
where $H^{{}^{\prime}}_{l}=H_{l}/P_{l}$, $W^{{}^{\prime}}_{l}=W_{l}/P_{l}$. The spatial goodness of each patch is computed as the mean squared activation (energy) within the patch region:
$g_{l,c}^{(i,j)}=\frac{1}{H^{{}^{\prime}}_{l}W^{{}^{\prime}}_{l}}\sum_{h=1}^{H^{{}^{\prime}}_{l}}\sum_{w=1}^{W^{{}^{\prime}}_{l}}\left(\mathbf{\tilde{Y}}_{l}^{\left(c,i,j,h,w\right)}\right)^{2}$ (2)
where $g_{l,c}^{(i,j)}$ denotes the spatial goodness of the $(i,j)$-th patch in channel $c$ at layer $l$. The spatial goodness values of all patches across all channels are concatenated to form the goodness vector $\mathbf{g}_{l}\in\mathbb{R}^{1\times\left(C_{l}\cdot P_{l}^{2}\right)}$.
Channel-Aware Patch Partitioning
As discussed above, spatial goodness extraction relies on partitioning feature maps into patches. However, applying a fixed partitioning scheme uniformly across all layers overlooks the varying levels of feature abstraction captured at different depths of a CNN (Zeiler and Fergus 2013). To address this, we propose a channel-aware patch partitioning strategy, in which a partitioning factor $P_{l}$ is adaptively scaled inversely with channel count $C_{l}$. Specifically, shallow layers tend to capture low-level, fine-grained features and thus benefit from finer spatial partitioning to produce sufficiently expressive goodness vectors for supervision. In contrast, deeper layers encode more abstract, high-level representations, where coarser spatial partitioning is sufficient to yield meaningful goodness vectors. This behavior is consistent with CNN architectures such as VGG (Simonyan and Zisserman 2015) and ResNet (He et al. 2015), where early layers typically have fewer channels and higher spatial resolutions, while deeper layers exhibit increased channel dimensionality and reduced spacial resolution. Consequently, it is intuitive to decrease $P_{l}$ as $C_{l}$ increases, as shown in Figure 2. We therefore compute $P_{l}$ as:
$P_{l}=\min\left(\max\left(1,\left\lfloor\frac{\alpha\cdot C_{L}}{C_{l}}\right\rfloor\right),H_{l},W_{l}\right)$ (3)
where $C_{L}$ is the channel count of the final convolutional layer, and $\alpha\geq 0$ is a hyperparameter that controls partitioning granularity. This formulation ensures that $P_{l}\geq 1$ , $H^{{}^{\prime}}_{l}\geq 1$ and $W^{{}^{\prime}}_{l}\geq 1$, while adaptively reducing the number of patches $P_{l}$ as channel dimensionality $C_{l}$ increases. A larger $\alpha$ results in finer partitioning with more patches, allowing the model to capture more detailed spatial goodness patterns. However, excessively large $\alpha$ may not yield further performance improvements. The effect of varying $\alpha$ will be further examined in the experimental section.
Random Projection for Supervision
The spatial goodness vectors obtained as described above are used for layer-wise supervision via classification. A common approach is to employ auxiliary classifiers (Belilovsky et al. 2019; Nøkland and Eidnes 2019); however, this introduces additional parameters and requires local backpropagation through the auxiliary branches, undermining the goal of achieving fully backpropagation-free training. To address this, we adopt a fixed random projection approach inspired by prior works (Huang et al. 2006; Huang and Aminifar 2025; Apolinario et al. 2025). Specifically, each layer’s goodness vector is projected into the $N$-dimensional class space via a linear transformation defined by a weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{\left(C_{l}\cdot P_{l}^{2}\right)\times N}$ and a bias vector $\mathbf{b}_{l}\in\mathbb{R}^{1\times N}$, both of which are randomly initialized at the beginning of training and remain fixed throughout. The generation procedure for these weight and bias parameters is as follows:
$\mathbf{W}_{l}\sim\mathcal{N}\left(0,\frac{1}{N}\right),\mathbf{b}_{l}\sim\mathcal{N}\left(0,\frac{1}{N}\right)$ (4)
The scaling factor $1/N$ for $\mathbf{W}_{l}$ follows the Johnson-Lindenstrauss lemma, ensuring that the $\ell_{2}$ norm of the projected vector remains close to that of the original goodness vector (Johnson and Lindenstrauss 1984). The bias vector $\mathbf{b}_{l}$ is similarly scaled by $1/N$, serving as a static offset that improves the logit distribution after projection. The classification logits $\mathbf{a}_{l}\in\mathbb{R}^{1\times N}$ are obtained from $\mathbf{g}_{l}$ via:
$\mathbf{a}_{l}=\mathbf{g}_{l}\mathbf{W}_{l}+\mathbf{b}_{l}$ (5)
which is then passed through a local cross-entropy loss $\mathcal{L}_{l}\left(\mathbf{a}_{l};t^{\ast}\right)$, computed against the ground-truth target $t^{\ast}$:
$\mathcal{L}_{l}\left(\mathbf{a}_{l};t^{\ast}\right)=-\mathbf{a}_{l}^{\left(t^{\ast}\right)}+\text{log}\sum_{n=1}^{N}{\text{exp}\left(\mathbf{a}_{l}^{\left(n\right)}\right)}$ (6)
The loss is subsequently used to compute gradients with respect to the convolutional parameters $\boldsymbol{\theta}_{l}$ of layer $l$:
$\nabla_{\theta_{l}}\mathcal{L}_{l}=\left(\frac{\partial\mathcal{L}_{l}}{\partial\mathbf{a}_{l}}\mathbf{W}_{l}^{\top}\right)\frac{\partial\mathbf{g}_{l}}{\partial\mathbf{Y}_{l}}\frac{\partial\mathbf{Y}_{l}}{\partial\boldsymbol{\theta}_{l}}$ (7)
This formulation allows the loss to directly produce gradients for the convolutional parameters, eliminating the need for local backpropagation through auxiliary trainable layers.
Training Framework
Overview
As illustrated in Figure 2, during one forward pass, each layer is trained independently using spatial goodness-based supervision. Specifically, the convolutional output at each layer is first passed through a ReLU nonlinearity. The resulting activated feature maps are then used to compute the layer’s spatial goodness and the corresponding loss, while simultaneously being forwarded to the subsequent operations—pooling (if present) and layer normalization (Ba et al. 2016). Critically, each layer’s output is detached before propagation to downstream layers, preventing gradient flow across layers and maintaining strict layer-wise independence.
Goodness-Preserving Pooling
Pooling is a standard downsampling operation in CNNs, used to reduce spatial resolution and computational complexity. However, to preserve spatial goodness information during downsampling, the pooling method must align with the energy-based formulation used to compute goodness. Standard approaches like max pooling and average pooling distort energy representation: the former tends to exaggerate energy by focusing on extreme values, while the latter may suppress it by smoothing activations. To maintain compatibility with our energy-based goodness metric, we adopt an approach inspired by Learned-Norm Pooling (Gulcehre et al. 2014) which generalizes pooling through the use of $L_{p}$ norm. Specifically, we set $p$ = 2, resulting in a root mean square (RMS) pooling. Given a localized region $\mathcal{R}\in\mathbb{R}^{H\times W}$ of a feature map, RMS pooling is defined as:
$\text{RMSPool}\left(\mathcal{R}\right)=\sqrt{\frac{1}{HW}\sum_{h}^{H}\sum_{w}^{W}\left(\mathcal{R}^{\left(h,w\right)}\right)^{2}}$ (8)
RMS pooling preserves the fundamental energy relationship, which directly aligns with our mean-squared-activation-based goodness computation, thus serves as an effective goodness-preserving downsampling mechanism.
Prediction Strategies
We introduce three strategies (Figure 2) for generating the final prediction during training and inference: (1) Last Layer Only prediction (Last Pred), (2) Cumulative Layer Fusion prediction (Fusion Pred), and (3) Best Layer Selection prediction (Best Pred).
In the Last Pred strategy, only the final convolutional layer’s features are passed through global average pooling (GAP), flattened, and fed into a single-layer linear classifier. During training, the resulting loss is used exclusively to update the classifier’s parameters, leaving earlier convolutional layers unaffected.
Given that each convolutional layer in our framework receives direct supervision from the ground-truth labels, each is expected to learn features that are independently useful for classification. To capitalize on this, the Fusion Pred strategy aggregates information from all convolutional layers except the first—excluded due to its low prediction accuracy. Specifically, we apply GAP and flattening to the feature maps of each selected layer, then concatenate the resulting representations. This combined vector is then passed to a single-layer linear classifier. Similar to the Last Pred strategy, only the classifier’s parameters are updated by the corresponding loss during training. This strategy effectively captures a broader range of features by leveraging the predictive capacity of multiple layers. However, it introduces additional classifier parameters and incurs memory overhead due to the need to store intermediate representations. This approach draws inspiration from the original FF method introduced in (Hinton 2022), but is specifically adapted to suit a convolutional architectures.
In the Best Pred strategy, each convolutional layer produces a classification output and associated prediction accuracy during training. The final prediction is then selected from the layer that achieves the highest validation accuracy. This approach offers two key advantages: (1) it eliminates the need for a final linear classifier, and (2) it enables early exit from inference when the most accurate prediction arises from a shallower layer, reducing computational cost.
Experiments
We evaluate our approach on five benchmark datasets: MNIST (Lecun et al. 1998), FashionMNIST (F-MNIST) (Xiao et al. 2017), CIFAR-10, CIFAR-100 (Krizhevsky and Hinton 2009), and ImageNet (ILSVRC 2012). To assess scalability, we include ImageNet’s 1.28 million training and 50,000 validation images across 1000 classes. For MNIST and F-MNIST, the original training sets are split into 50,000 training and 10,000 validation samples, and models are trained for 100 epochs. For CIFAR-10 and CIFAR-100, we use 45,000 training and 5,000 validation samples, with training conducted over 400 epochs. For all datasets except ImageNet, we adopt the AdamW optimizer with a weight decay of 0.001, a batch size of 128 and an initial learning rate of 0.0002, which is decayed to 0.00001 using a cosine annealing scheduler. For ImageNet, we use the SGD optimizer with momentum 0.9 and weight decay 0.0001. The batch size is 512, and the learning rate is scheduled from 0.05 to 0.0005 via cosine annealing. Models are trained for 45 epochs. Training is performed on 2 NVIDIA A100 GPUs (80 GB) and 1 Intel(R) Xeon(R) Gold 6338 CPU. We report the mean accuracy and standard deviation averaged over three independent runs. In each run, the model checkpoint with the highest validation accuracy is selected for final evaluation on the test set. Additional details, including model architectures, data augmentation strategies and so on, are provided in the supplementary materials.
Ablation Study
Figure 3: Layer-wise prediction accuracy (layers 2-7) of VGG8 trained with ASGE on CIFAR-10. First and final layers excluded. Left: training. Right: validation.
ASGE ($\alpha$) Model Partition ($P_{l}$) MNIST F-MNIST CIFAR-10 CIFAR-100
0.0 VGG8 {1, 1, 1} 99.61 $\pm$ 0.01 92.88 $\pm$ 0.08 88.27 $\pm$ 0.20 61.69 $\pm$ 0.23
0.5 VGG8 {2, 1, 1} 99.64 $\pm$ 0.03 92.79 $\pm$ 0.10 89.20 $\pm$ 0.07 62.31 $\pm$ 0.29
1.0 VGG8 {4, 2, 1} 99.65 $\pm$ 0.03 93.41 $\pm$ 0.14 90.50 $\pm$ 0.08 65.34 $\pm$ 0.31
1.5 VGG8 {6, 3, 1} 99.63 $\pm$ 0.01 93.36 $\pm$ 0.07 90.22 $\pm$ 0.13 64.84 $\pm$ 0.35
2.0 VGG8 {8, 4, 2} 99.57 $\pm$ 0.04 93.12 $\pm$ 0.10 86.91 $\pm$ 0.22 60.40 $\pm$ 0.23
Table 1: Test accuracy (%) of ASGE under different values of $\alpha$ across four datasets. VGG8 contains layers with $C_{l}\in$ {128, 256, 512} and Partition row shows the corresponding $P_{l}$ assigned to each $C_{l}$ according to formula (3).
Figure 4: Spatial goodness distributions on a logarithmic scale before (G-Origin) and after different pooling (G-RMS, G-Avg, G-Max) in a layer that has pooling. Each row represents a different channel configuration: C = 512 (top) and C = 256 (bottom).
ASGE (Pooling) Model MNIST F-MNIST CIFAR-10 CIFAR-100
Average VGG8 99.62 $\pm$ 0.03 93.18 $\pm$ 0.10 90.35 $\pm$ 0.15 64.96 $\pm$ 0.16
Max VGG8 99.61 $\pm$ 0.03 93.07 $\pm$ 0.16 90.46 $\pm$ 0.18 64.41 $\pm$ 0.11
RMS VGG8 99.65 $\pm$ 0.03 93.41 $\pm$ 0.14 90.50 $\pm$ 0.08 65.34 $\pm$ 0.31
Table 2: Test accuracy (%) of ASGE with different pooling operations.
Prediction Strategy Model Test Accuracy (%) Activation Memory (MB) Classifier Params ($\times$104) Best Layer
BP (repr.) VGG8 69.11 $\pm$ 0.31 150 5.13 -
ASGE Fusion Pred VGG8 65.34 $\pm$ 0.31 1.25 25.66 -
ASGE Last Pred VGG8 63.31 $\pm$ 0.20 0.25 5.13 -
ASGE Best Pred VGG8 63.60 $\pm$ 0.44 0 0 7
Table 3: CIFAR-100: ASGE prediction strategies vs. BP. Activation Memory: calculated using a batch of 128. Best Layer: the layer with the highest validation accuracy.
ImageNet
Method Model MNIST F-MNIST CIFAR-10 CIFAR-100 Top-1 Top-5
[…]
FA MLP8 98.74 - 58.03 - 6.92 17.46
DTP VGG6 98.93 $\pm$ 0.04 90.91 $\pm$ 0.17 85.33 $\pm$ 0.32 - 1.66 5.44
DFA CNN5 98.98 $\pm$ 0.05 - 73.10 $\pm$ 0.50 41.00 $\pm$ 0.30 6.20 -
SoftHebb SoftHebb 99.35 $\pm$ 0.03 - 80.31 $\pm$ 0.14 56.00 27.30 -
rLRA ResNet18 98.18 88.13 93.58 - 69.22 87.96
PEPITA CNN2 98.29 $\pm$ 0.13 - 56.33 $\pm$ 1.35 27.56 $\pm$ 0.60 - -
$F$ MLP4 98.69 - 59.00 - - -
CaFo CNN3 98.80 - 67.43 40.76 - -
SymBa MLP3 98.58 - 59.09 29.28 - -
[…]
DeeperForward ResNet18 99.63 $\pm$ 0.04 93.13 $\pm$ 0.13 88.72 $\pm$ 0.17 53.09 $\pm$ 0.79 - -
ASGE (ours) VGG8 99.65 $\pm$ 0.03 93.41 $\pm$ 0.14 90.50 $\pm$ 0.08 65.34 $\pm$ 0.31 23.04 43.08
ASGE (ours) VGG11 99.59 $\pm$ 0.06 92.98 $\pm$ 0.24 90.62 $\pm$ 0.17 65.42 $\pm$ 0.16 26.21 47.49
ASGE(ours) ResNet18-CHx4 - - - - 51.58 75.23
[…]
Effect of ASGE
Figure 1 demonstrated a core limitation of FF training: a progressive degradation of layer-wise prediction accuracy with increasing network depth in CNNs. To assess whether ASGE mitigates this issue, Figure 3 presents the layer-wise accuracy of our VGG8 architecture, excluding the first layer—due to its negligible predictive performance—and the final linear classifier. The results reveal a fundamental reversal: from layers 2–7, prediction accuracy consistently improves across both the training and validation sets. Notably, the inter-layer accuracy gain diminishes with depth, indicating stabilization of increasingly refined feature representations. Overall, these findings confirm that ASGE successfully counteracts the FF’s representational degradation, enabling effective information accumulation across layers for scalable deep network learning.
Effect of Partitioning Factor $\alpha$
To examine the impact of the hyperparameter $\alpha$ on ASGE performance, we evaluate five representative values: 0, 0.5, 1, 1.5, and 2. All other training configurations are held constant, including the use of RMS pooling and the Fusion Pred strategy for final prediction. The results are summarized in Table 1. Across all datasets, a consistent trend emerges: test accuracy improves as $\alpha$ increases from 0 to 1, but begins to decline when $\alpha$ exceeds 1. The best performance is achieved at $\alpha$ = 1, yielding test accuracies of 99.65% on MNIST, 93.41% on F-MNIST, 90.50% on CIFAR-10, and 65.34% on CIFAR-100. We interpret this behavior as follows: increasing $\alpha$ results in more spatial partitions per layer, allowing each layer to capture richer and more fine-grained spatial goodness information, which enhances performance. However, when $\alpha$ grows too large, the surplus patches introduces redundant information and noise, causing potential overfitting and degraded generalization. Thus, $\alpha$ = 1 provides an effective balance between spatial granularity and noise, enabling optimal performance.
Effect of RMS Pooling
Figure 4 visualizes the impact of different pooling operations on spatial goodness distributions. The first two columns show the distributions from the original representation (G-Origin) and after applying RMS pooling (G-RMS). The subsequent two columns compare G-Origin with the distributions obtained after average pooling (G-Avg) and max pooling (G-Max), respectively. It is evident that G-RMS closely mirrors G-Origin, preserving both the tail behavior and the overall distribution structure. In contrast, G-Avg distorts the original distribution by suppressing the tail and shifting toward lower goodness scores. G-Max, on the other hand, exaggerates the tail, increasing the proportion of higher goodness values. This indicates that RMS pooling more faithfully preserves spatial goodness information during downsampling, whereas the other two pooling methods introduce distributional distortions. Furthermore, we train a VGG8 using each pooling method with $\alpha$ = 1 and the Fusion Pred strategy for final prediction. As shown in Table 2, ASGE with RMS pooling consistently outperforms average and max pooling across all datasets. By more efficiently preserving goodness information, RMS pooling aligns better with the core principles of ASGE and enhances the model’s representational capacity.
Performance of Prediction Strategies
[…]
Comparison with Other Work
We benchmark ASGE against standard BP and diverse BP-free approaches, including both FF-based and non-FF methods. Among non-FF methods, we include FA (Bartunov et al. 2018), Difference Target Propagation (DTP) (Ernoult et al. 2022; Bartunov et al. 2018), DFA (Nøkland 2016; Crafton et al. 2019), SoftHebb (Journé et al. 2023), rLRA (Ororbia et al. 2023) and PEPITA (Dellaferrera and Kreiman 2023)). For FF-based methods, we evaluate the original FF (Hinton 2022), CaFo (Zhao et al. 2023), SymBa (Lee and Song 2023), CwC (Papachristodoulou et al. 2024) and DeeperForward (Sun et al. 2025). All experiments for ASGE are conducted using VGG8 and VGG11 with $\alpha$ = 1, RMS pooling, and Fusion Pred for final prediction.
Table 4 shows that ASGE consistently outperforms prior FF-based methods across MNIST, F-MNIST, CIFAR-10, and CIFAR-100, and surpasses most non-FF BP-free methods. The only exception is rLRA, which achieves a 2.96% higher accuracy on CIFAR-10. Notably, on the more challenging CIFAR-100 dataset, ASGE shows a significant advantage, exceeding all previous FF-based methods by at least 12.33%. Compared to BP, ASGE delivers slightly lower but still competitive results, with accuracy gaps of just 0.88% on F-MNIST, 1.18% on CIFAR-10, and 3.77% on CIFAR-100.
To our knowledge, ASGE is the first FF-based method to successfully scale to ImageNet, achieving Top-1 and Top-5 accuracies of 26.21% and 47.49%, setting a new benchmark for FF-based training. ASGE outperforms non-FF BP-free methods like FA, DTP, and DFA, though it trails SoftHebb, rLRA, and standard BP. However, rLRA requires a full forward pass before updates, increasing memory usage, while SoftHebb is limited to specific network architectures. ASGE circumvents these constraints with fully layer-wise training and flexible integration with convolutional architectures.
Conclusion
This work introduces ASGE, a novel FF-based BP-free training framework for CNNs that overcomes key limitations of prior FF methods through channel-aware spatial goodness supervision, enhancing both representational capacity and scalability to large-scale datasets. Experiments demonstrate that ASGE consistently outperforms all FF-based methods across 5 image classification benchmarks, and further narrows the performance gap with BP. Notably, it marks the first successful application of FF-based training to ImageNet, demonstrating its potential as a scalable and effective BP-free paradigm for modern CNNs.
Limitations
Despite its strengths, ASGE has certain limitations. As a greedy, layer-wise approach, it generally requires longer training time to converge than end-to-end BP. Currently, ASGE is only applied to CNNs, and its extension to other architectures such as Transformers or RNNs remains to be explored. Additionally, while ASGE performs competitively across various datasets, it still trails some non-FF BP-free methods on ImageNet, indicating room for future improvement.